my daughter is doing math quiz and doesnt under stand how to find the answer to this problem... find the sum of the first 100 odd numbers?
Answer by ankor@dixie-net.com(12701)

find the sum of the first 100 odd numbers?
:
You know there are 100 odd numbers between 0 and 200
Each can be paired, so there are 50 pairs:
:
199 + 1 = 200
197 + 3 = 200
195 + 5 = 200
193 + 7 = 200
///////////
101 + 99 = 200
:
Wouldn't lt just be, 50 * 200 = 10000?